### v1

@@ 5 Mapping Drivers to Barriers
In this section, we will map the drivers of reproducibility to the barriers in the form of a Drivers-Barriers Matrix. This will be based on the definition and categorization of reproducibility as a foundation (Section 2), and the identification of the major barriers (Section 3) and drivers (Section 4) of ML reproducibility. The resulting Drivers-Barriers-Matrix is depicted in Figure 2 and categorizes the barriers into the four different types of ML reproducibility, i.e., R1 Description, R2 Code, R3 Data, R4 Experiment [25], and drivers into technology-driven, procedural, and related to awareness and education.
@@
The major drivers for completeness and report quality are guidelines and checklists. Guidelines provide best practices to adopt in order to achieve reproducible ML research. Furthermore, many checklists exist that comprehensively state the different pitfalls and provide information on how they can be avoided. Researchers can use them to ensure their research meets the desired standards. Furthermore, some of these checklists and guidelines are enforced by journals, such that research will only be published if certain criteria are met.
In comparison, spin practices are not as easily identifiable. In this case, the discussion within the research community centers around removing the incentives for inflating research results. A particularly effective driver for this is preregistration as an example for publication policies and initiatives, where researchers submit research proposals before conducting an experiment. If accepted, the research will be published, regardless of the outcome, thereby minimizing spin practices.
@@
Code sharing is essential to reproducibility, which makes limited access to code a significant barrier in the field. However, it is often neglected as the process is not trivial. To make shared code useful to the scientific community, it is necessary to share, in addition to the source code, the information about the entire software setup and dependencies, including software versions and hardware configurations. To assist with this, researchers may consider running their code in hosting services or virtualization environment, which we identified as drivers for code sharing. Both have similar advantages, i.e., they can easily be shared and made public for other researchers to use. As a consequence, it will give reproducers immediate access to code, including the complete configuration setup, such as dependencies and versions. Hosting services are the quicker and easier way of achieving this; however, they may be subject to different resource limits. Virtualization (e.g., VMs or containers) is more difficult to set up but offers more flexibility and is not externally (e.g., by a provider) restricted in capabilities and resources. Furthermore, tools and platforms can be drivers for reproducibility. A lot of ML tools provide capabilities for code versioning or other features, which are key to reproducibility. One example is dToolAI [31], which automatically logs the supplemental information of the code, i.e., important metadata, hyperparameters, and more, which are essential to the reproducibility of type R2 Code.
@@
Bias is another methodological error, leading to irreproducible results. This is because the biased data usually does not generalize well to problems outside the experiment setup. Biases have been an important source of concern, e.g., in Biomedical fields. Effects thereof can again be minimized using standardized datasets and evaluation or specific guidelines and checklists, e.g., ROBUST-ML [2].
@@
If an ML experiment is shared entirely and code and data are available, i.e., reproducibility type R4 Experiment, there are still three barriers, which can lead to irreproducible results. Inherent non-determinism arises from the different sources of randomness in ML, and makes it difficult to achieve repeatable results on the same machine. There are, however, methods to manage the sources of randomness, such as fixed random seeds and deterministic implementations, while comprehensively mitigating all sources of randomness is still a very challenging endeavor.
Another barrier is described as environmental differences, which has two main issues associated with it, i.e., software differences and hardware differences. Both types of differences can be avoided by using either hosting services or virtualization; constraints can be assumed to be similar to the barrier of limited access to code.
Limited access to computational resources constitutes another barrier to ML reproducibility identified in this work. The issue is particularly problematic for research on large language models because of their need for extensive computational resources in training and reproduction. Hosting services offer a solution, providing access to pre-trained models and allowing researchers to directly access and run respective models on-site.
@@ 6 Conclusion
In this paper, we examined the barriers and drivers associated with the four types of ML reproducibility as outlined by Gundersen et al. (Description, Data, Code, and Experiment) [25], specifically in the contexts of Computer Science and Biomedical research. We then synthesized this into a Drivers-Barriers-Matrix to summarize and illustrate which drivers are feasible solutions for the various barriers. We observe that the barriers to machine learning reproducibility can be addressed through three primary drivers: technology-driven solutions, procedural improvements, and enhanced awareness and education. It is important to highlight that, in theory, awareness and education can complement the other drivers and serve as a foundational step toward overcoming any reproducibility-related challenge. Further, the rise of AutoML tools for ML development and ML tasks performed by domain experts (potentially) not having in-depth computer or data science knowledge could become another barrier for reproducibility. However, we are confident that our findings will also help in this regard and that the topic likely requires collaborative efforts over the coming years.
We hope that this provides practical guidance and orientation for ML researchers, and clarifies the current state of play. Of course, in such a dynamic and fast-paced research area, this only opens a series of further questions and avenues for exploration. We recommend further investigation of the various issues and potential solutions laid out here. We would also encourage further investigation into the potential role of platforms [29] or foundation models [34] in further exacerbating or alleviating these issues.

### v2

@@ 5 Mapping Drivers to Barriers
In this section, we map the drivers of reproducibility to the barriers in the form of a Drivers-Barriers Matrix. This will be based on the definition and categorization of reproducibility as a foundation (Section 2), and the identification of the major barriers (Section 3) and drivers (Section 4) of ML reproducibility. The resulting Drivers-Barriers-Matrix is depicted in Figure 2 and categorizes the barriers into the four different types of ML reproducibility, i.e., R1 Description, R2 Code, R3 Data, R4 Experiment [26], and drivers into technology-driven drivers, procedural drivers, and drivers related to awareness and education.
@@
The major drivers for completeness and report quality are guidelines and checklists. Guidelines provide best practices to adopt in order to achieve reproducible ML research. Furthermore, many checklists exist that comprehensively state the different pitfalls and provide information on how they can be avoided. Researchers can use them to ensure their research meets the desired standards. Furthermore, some of these checklists and guidelines are enforced by journals, such that research will only be published if certain criteria are met. In comparison, spin practices are not as easily identifiable. In this case, the discussion within the research community centers around removing the incentives for inflating research results. A particularly effective driver for this is preregistration as an example for publication policies and initiatives, where researchers submit research proposals before conducting an experiment. If accepted, the research will be published, regardless of the outcome, thereby minimizing spin practices.
@@
Code sharing is essential to reproducibility, which makes limited access to code a significant barrier in the field. However, it is often neglected as the process is not trivial. To make shared code useful to the scientific community, it is necessary to share, in addition to the source code, the information about the entire software setup and dependencies, including software versions and hardware configurations. To assist with this, researchers may consider running their code in hosting services or virtualization environments, which we identified as drivers for code sharing. Both have similar advantages, i.e., they can easily be shared and made public for other researchers to use. As a consequence, it will give reproducers immediate access to code, including the complete configuration setup, such as dependencies and versions. Hosting services are the quicker and easier way of achieving this; however, they may be subject to different resource limits. Virtualization (e.g., VMs or containers) is more difficult to set up but offers more flexibility and is not externally (e.g., by a provider) restricted in capabilities and resources. Furthermore, tools and platforms can be drivers for reproducibility. A lot of ML tools provide capabilities for code versioning or other features, which are key to reproducibility. One example is dToolAI [32], which automatically logs the supplemental information of the code, i.e., metadata, hyperparameters, and more, which are essential for ML reproducibility.
@@
Bias is another methodological error, leading to irreproducible results. This is because the biased data usually does not generalize well to problems outside the experimental setup of a specific ML study. Bias has been an important source of concern, e.g., in Biomedical fields. Effects thereof can again be minimized using standardized datasets and evaluation or specific guidelines and checklists, e.g., ROBUST-ML [2].
@@
If an ML experiment is shared entirely and code and data are available, i.e., reproducibility type R4 Experiment, there are still three barriers, which can lead to irreproducible results. Inherent non-determinism arises from the different sources of randomness in ML, and makes it difficult to achieve repeatable results, even on the same machine. There are, however, methods to manage the sources of randomness, such as fixed random seeds and deterministic implementations, while comprehensively mitigating all sources of randomness is still a very challenging endeavor.
Another barrier is described as environmental differences, which has two main issues associated with it, i.e., software differences and hardware differences. Both types of differences can be avoided by using either hosting services or virtualization; constraints can be assumed to be similar to the barrier of limited access to code. Limited access to computational resources constitutes another barrier to ML reproducibility identified in this work. The issue is particularly problematic for research using LLMs because of their need for extensive computational resources in training and reproduction. Hosting services offer a solution, providing access to pre-trained models and allowing researchers to directly access and run respective models on-site.
@@ 6 Conclusion
In this paper, we examined the barriers and drivers associated with the four types of ML reproducibility as outlined by Gundersen et al. (description, data, code, and experiment) [26], specifically in the contexts of Computer Science and Biomedical research. We then synthesized this into a Drivers-Barriers-Matrix to summarize and illustrate which drivers are feasible solutions for the various barriers. We observe that the barriers to ML reproducibility can be addressed through three primary drivers: technology-driven solutions, procedural improvements, and enhanced awareness and education. It is important to highlight that, in theory, awareness and education can complement the other drivers and serve as a foundational step toward overcoming any reproducibility-related challenge. Further, the rise of AutoML tools for ML development and ML tasks performed by domain experts, (potentially) not having in-depth computer or data science knowledge, could become another barrier for reproducibility. However, we are confident that our findings will also help in this regard and that the topic likely requires collaborative efforts over the coming years.
We hope that our paper provides practical guidance and orientation for researchers employing ML, and clarifies the current state of play. Of course, in such a dynamic and fast-paced research area, this only opens a series of further questions and avenues for exploration. We recommend further investigation of the various issues and potential solutions laid out here. We would also encourage further investigation into the potential role of platforms [30] or foundation models [35] in further exacerbating or alleviating these issues.